OpenStudy (anonymous):

Which circles have an area of 16pi in2? Choose all answers that are correct. A. a circle with a radius of 2 in. B. a circle with a radius of 4 in. C. a circle with a diameter of 8 in. D. a circle with a diameter of 16 in.

jhonyy9 (jhonyy9):

do you know formula for area of a circle ?

OpenStudy (aihberkhan):

To find the area of a circle you have to solve \(\bf \large \pi r^2\)

OpenStudy (aihberkhan):

Now \(\bf r \) is the radius.

OpenStudy (anonymous):

@jhonyy9 yes its area= pixradious squared

OpenStudy (aihberkhan):

In the answer choices they have given you the radius. So plug in each radius into the formula and find which one has the answer of 16 pi in squared

OpenStudy (anonymous):

@AihberKhan can u help me through it?

OpenStudy (aihberkhan):

Sure! :) @parisjackson

OpenStudy (anonymous):

got my paper, lets go

OpenStudy (aihberkhan):

Haha Okay xD So the 1st one, when you plug in 2 for r it should look like \(2 \pi in^2\)

OpenStudy (aihberkhan):

And that is NOT what we want. So we can eliminate A.

OpenStudy (anonymous):

got it

OpenStudy (aihberkhan):

Good! :) Now the next one... for radius we have to plug in \(4\). Why don't you do this... when you plug in 4 for \(r\) what would you get? @parisjackson

OpenStudy (anonymous):

4piinsquared

OpenStudy (aihberkhan):

Right! :) Now that is NOT what we want... so we can eliminate B... Do you see the pattern? All we need is the number 16.. which answer has that number? So what do you think the answer is? Do you know now? :) @parisjackson

OpenStudy (anonymous):

its d

OpenStudy (aihberkhan):

Thats what I thought... but now I am looking at it and it says 16 Diameter... not radius... None of them have 16 for radius, so there is no answer... :P How about we ask someone else to \(\bf \text {double check}\) :)

OpenStudy (aihberkhan):

@UsukiDoll Can you double check... we are a bit confused... :)

OpenStudy (aihberkhan):

@jim_thompson5910 @jhonyy9 Can you double check... we are a bit confused... :)

OpenStudy (aihberkhan):

@TheSmartOne @pooja195 Can you double check... we are a bit confused... :)

jimthompson5910 (jim_thompson5910):

when you plug in the radius of 2, it should be 4pi square inches notice how 2 is squared to get 4

OpenStudy (anonymous):

x2

OpenStudy (aihberkhan):

Right but don't we need an answer that is squared?

jimthompson5910 (jim_thompson5910):

A = pi*r^2 A = pi*2^2 A = pi*4 A = 4pi

OpenStudy (anonymous):

yeah, it has to be squared

jimthompson5910 (jim_thompson5910):

`Right but don't we need an answer that is squared? ` you mean the units are squared?

OpenStudy (aihberkhan):

Oh wait I get it now! Yeah follow what @jim_thompson5910 is saying :) Sorry I couldn't help you better.. :) @parisjackson

jimthompson5910 (jim_thompson5910):

think of it like this (2 inches) * (2 inches) = (2*2)*(inches*inches) = 4*(inches)^2 only the "inches" are squared at the end

jimthompson5910 (jim_thompson5910):

so a circle with radius 2 inches leads to an area of 4pi square inches we want 16pi square inches so r = 2 is NOT the answer

jimthompson5910 (jim_thompson5910):

there's only one value of r that works try r = 4 now

OpenStudy (anonymous):

8???

jimthompson5910 (jim_thompson5910):

4 squared = 4*4 = 16

jimthompson5910 (jim_thompson5910):

`4 squared` is not the same as `4 times 2`

OpenStudy (anonymous):

yeah yeah yeah i get it

jimthompson5910 (jim_thompson5910):

you'll also use the idea that diameter = 2*(radius)

jimthompson5910 (jim_thompson5910):

for example, if the radius was 5 then the diameter is 10 diameter = 2*(radius) diameter = 2*(5) diameter = 10

OpenStudy (anonymous):

omg thanks

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

so its b?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

B) because r = 4 inches leads to an area of 16pi square inches AND C) because a diameter of 8 inches leads to a radius of r = 4 inches (use the formula connecting diameter and radius)

jimthompson5910 (jim_thompson5910):

`a circle with a radius of 4 in.` and `a circle with a diameter of 8 in.` are the same circle (with the same area)

OpenStudy (anonymous):

@jim_thompson5910 i have another problem almost the same. can i get some help?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

ok give me a second

OpenStudy (anonymous):

What is the exact area of a circle with a diameter of ft? A. square feet B. square feet C. square feet D. square feet

OpenStudy (anonymous):

didnt come out right

OpenStudy (anonymous):